### v1

# Configurable DOA Estimation using Incremental Learning

Abstract
This study introduces a progressive neural network (PNN) model for direction of arrival (DOA) estimation, DOA-PNN, addressing the challenge due to catastrophic forgetting in adapting dynamic acoustic environments. While traditional methods such as GCC, MUSIC, and SRP-PHAT are effective in static settings, they perform worse in noisy, reverberant conditions. Deep learning models, particularly CNNs, offer improvements but struggle with a mismatch configuration between the training and inference phases. The proposed DOA-PNN overcomes these limitations by incorporating task incremental learning of continual learning, allowing for adaptation across varying acoustic scenarios with less forgetting of previously learned knowledge. Featuring task-specific sub-networks and a scaling mechanism, DOA-PNN efficiently manages parameter growth, ensuring high performance across incremental microphone configurations. We study DOA-PNN on a simulated data under various mic distance based microphone settings. The studies reveal its capability to maintain performance with minimal parameter increase, presenting an efficient solution for DOA estimation.
Index Terms—  DOA Estimation, Sound Source Localization, Incremental Learning, Configurable Model
1 Introduction
Sound source localization (SSL) [1] aims to identify the origins of sound sources automatically by analyzing signals captured through an array of microphones. At the cornerstone of SSL lies the direction of arrival (DOA) estimation, a process crucial for calculating the angles at which sounds reach the microphone array. This spatial information is necessary for a broad range of speech-related applications, including but not limited to, speech separation [2], speaker extraction [3], speech recognition [4], and speech enhancement [5]. These applications rely on the precise localization of sound sources to significantly improve their performance and functionality. Consequently, the development of more sophisticated DOA estimation techniques has become a key focus within the research community to achieve greater robustness and adaptability in real-world settings.
Classical signal processing techniques like generalized cross-correlation (GCC) [6], multiple signal classification (MUSIC) [7], and steered response power (SRP) based methods, particularly SRP-PHAT [8], which is obtained from the SRP by applying phase transform (PHAT) whitening, have been foundational due to their simplicity and versatility. They are available to adapt to different microphone array configurations. However, their effectiveness decreases in noisy and reverberant environments, which leads to exploration of deep learning solutions.
A significant milestone in this exploration is the adoption of deep learning architectures such as convolutional neural networks (CNN) [9, 10, 11, 12, 13]. However, deep learning models used for DOA estimation face a unique challenge: the misalignment between training and testing environments, especially concerning the microphone array’s configuration. Such mismatches, particularly changes in microphone inter-distances, precipitate a decline in model efficacy, necessitating re-training for new configurations. Re-training is not always feasible and requires extra time or computation resources. This highlights an essential area for ongoing research and innovation of deep learning-based DOA estimation, to enhance model flexibility and accuracy across diverse acoustic settings.
Recent advancements in deep DOA estimation [14] tackle array configuration mismatching by framing DOA as a classification problem within each time-frequency (TF) bin, using high-resolution images from narrow-band SRP-PHAT for input. With such high-level localization cues, relatively simple neural networks can be used. Although these methods adapt to the unseen array setups relying on specific spatial features, they do not address the issue when incrementally adjusting to new acoustic configurations. Along the diverse configurations, one fixed method faces forgetting because the new knowledge overrides the existing knowledge.
Various research recently incorporated continual learning (CL) [15] to continuously learn new knowledge while retaining and reusing previously learned knowledge in keyword spotting [16], audio classification [17], and sound event detection. Task Incremental Learning (TIL)  [18] is a subset of CL methods. The TIL approach enables the model to perform new tasks efficiently with the same number of classes in each task without losing prior performance. Inspired by [19, 20], we proposed to use TIL to gradually learn new information for DOA estimation using simple, widely used features like the short-time Fourier transform (STFT), all while keeping the valuable knowledge it has already gained from the existing acoustic environments and configurations.
In this research, we introduce DOA-PNN, a novel progressive neural network based on the TIL method for DOA estimation, crafted to solve the issue of adaptation from source to target configurations. The core of DOA-PNN’s architecture is task-specific sub-networks, each engineered to memorize knowledge from previous configurations, enabling smooth transitions across diverse acoustic environments while maintaining accuracy on previous tasks. A scaling mechanism is integrated into these sub-networks to limit parameter growth by reducing the extra layers. Contrary to traditional finetuning methods, DOA-PNN utilizes a collection of pre-trained sub-networks, establishing lateral connections that extract the crucial features for new tasks, thereby enriching feature presentation. To the best of our knowledge, we are the first to apply continual learning techniques to adapt DOA estimation to various acoustic configurations. Demonstrated through extensive testing on a simulated dataset for DOA estimation, DOA-PNN exhibits competitive performance on unfamiliar configurations with less parameter increase and no reliance on previous data buffers.
Figure 1: (a) Architecture of SSNet (b) Proposed DOA-PNN framework.
2 Proposed Method
2.1 Continual learning for DOA estimation
In this work, applying continual learning to DOA estimation involves receiving a series of $T$ distinct DOA estimation tasks (under different acoustic configurations) in sequence and optimizing the performance across all tasks with less negative impact of catastrophic forgetting. For each task $\tau_{t}$, with $t\leq T$, we handle a set of training data consisting of pairs $(x_{t},y_{t})$, where $x_{t}$ represents the acoustic signals and $y_{t}$ are the corresponding DOA labels, sampled from a distribution $D^{T}$.
Our goal is to minimize the cumulative loss $L_{tot}$, which aggregates the sub-losses $L_{doa}$ for all $T$ tasks, formulated as:
$L_{tot}=\sum_{t=0}^{T}E_{(x_{t},y_{t})\sim D^{T}}[L_{doa}(F_{t}(x_{t};\mu_{t}),y_{t})]$ (1)
where $L_{doa}$ represents a classification loss function, such as mean squared error (MSE) loss, and $F_{t}(x_{t};\mu_{t})$ is the DOA estimation model parameterized by $\mu_{t}$. During the training phase, we aim to find the optimal parameters $\mu_{t}^{\prime}$ that perform well across all $T$ tasks. This is challenging since parameters learned from previous task $\tau_{t}$ are prone to be overwritten when learning the subsequent task $\tau_{t+1}$, leading to catastrophic forgetting.
2.2 Network backbone for DOA estimation
We adopt a spatial spectrum prediction module using the spatial spectrum network (SSNet) [21, 22, 23] for DOA estimation. We consider 2 microphone linear arrays as an instance in this section, although our proposed method can be employed for any number of microphone positions. This system learns complex relationships from audio data to predict, where the sound could be coming from at every degree in a half-circle of the linear microphone array. To perform this, we combine sound segments from different microphones and transform them into complex STFT. We consider the real as well as the imaginary part of the complex STFT instead of the phase and power and then combine this data to form a feature representation $v$. This $v$ feature is then used as the input to SSNet for predicting the spatial spectrum $\hat{p}$. The network structure of SSNet is shown in Figure 1 (a) that depicts SSNet includes layers to simplify and analyze sound features. It uses two convolutional layers to reduce frequency details and five residual blocks to pick up intricate sound features. Then, it has a layer that focuses on the DOA estimation and rearranges the sound information in preparation for the final step. The result spatial spectrum is a 181-point map.
Following the work in [23], we use a probability-based method to determine the likelihood of a sound coming from any of the 180 directions. This part of SSNet assigns a probability to each direction based on how closely it matches the ground truth DOA. Specifically, each element of the encoded $181$-dimensional vector $p(\theta_{i})$ is assigned to a particular azimuth direction $\theta_{i}\in\{0^{\circ},1^{\circ},\dots,180^{\circ}\}$. The values of the vector follow a Gaussian distribution that maximizes at the ground truth DOA, which is defined as follows:
$p(\theta_{i})=\begin{cases}\exp\left(-\frac{d(\theta_{i},\theta_{y})^{2}}{2\sigma^{2}}\right),&\text{if }|\Theta|=0\\ 1,&\text{if }|\Theta|=1\end{cases}$ (2)
where $|\Theta|$ is the number of sources, $\theta^{\prime}$ is the ground truth DOA of one source, $\sigma$ is a predefined constant that controls the width of the Gaussian function and $d(\dots,\dots)$ denotes the angular distance. We then train SSNet by computing the MSE loss between the segment-wise output spatial spectrum from the model and the ground truth.
2.3 Proposed DOA-PNN
Drawing on the concept of progressive neural networks, we introduce a novel progressive continual learning approach specifically designed for deep DOA estimation, depicted in Figure 1(b). Our method, termed DOA-PNN, is composed of two principal elements: a task-specific sub-network creator for individual tasks and a lateral connection set bridging a suite of sub-networks. To manage the model size, we’ve introduced a gap-aware layer scaling mechanism in DOA-PNN, ensuring the parameter count remains controlled.
A task-specific sub-network creator is designed to generate the $t^{{\text{th}}}$ sub-network for each new task $\tau_{t}$. When a new acoustic configuration DOA estimation $\tau_{t}$ comes, the creator adds a sub-network with random initialization which is only the feature extractor part of SSNet. All sub-networks share the common memory which is the classification head of SSNet for a fixed class number 181.
A lateral connection set serve a pivotal role. When the $t^{{\text{th}}}$ sub-network processes input data, it not only forwards its output to the classification layer but also receives output features from all previously established sub-networks. To integrate these multiple outputs effectively, each sub-network is equipped with an adapter—a 1$\times$1 convolutional layer—tailored to its current context. This adapter transforms the combined features into the final output, thereby enriching the feature used for classification with the accumulated knowledge from the predecessors.
A gap-aware layer scaling mechanism adjusts the complexity of the SSNet feature extractor’s residual blocks based on a consistent parameter reflecting the minimum angle error tolerance. For instance, if the smallest acceptable angle error is 5 degrees, we streamline the residual blocks from five down to one. This scaling ensures the network remains efficient and tailored to the specific precision requirements of the DOA estimation task.
In this architecture, for each new learning task labeled as $\tau_{t}$, we compute the real and imaginary parts of the complex STFT of audio segments to serve as input data. Concurrently, the DOA-PNN model preserves the integrity of the previously trained $t-1$ sub-networks by freezing their parameters, preventing interference with newly acquired knowledge. The creator then creates a fresh $t^{{\text{th}}}$ sub-network dedicated to the current learning task, leveraging both the STFT features and the task metadata—such as task id—to discern previously unidentified acoustic configurations. What sets the $t^{{\text{th}}}$ sub-network apart is its ability to access the shared memory of its predecessors through lateral connections, integrating all previous knowledge. This structure not only facilitates effective knowledge transfer across various tasks, but also promotes faster convergence during the training process. When it comes to inference application, the DOA-PNN model selects the appropriate $t^{{\text{th}}}$ sub-network based on the specific task at hand for accurate and reliable evaluation. This approach ensures that our model remains adaptable and quick to respond in diverse acoustic settings, showcasing the potential of progressive continual learning strategy in the real-world environments of DOA estimation.
3 Experiments
Table 1: The parameter settings of simulated audio
Room Size [L, W, H] Absorption Tmax Num_images
[10,8,5], [5,4,3], [4,2,2] 0.2-0.8 0.2-0.6 [10,10,10]
3.1 Dataset
We use the LibriSpeech [24] train-clean-100 subset, which comprises high-fidelity, single-channel speech recordings sampled at 16 kHz. It includes a diverse set of English speakers, selected for uniform audio segment lengths, ensuring homogeneity in data processing. From this dataset, we have curated a selection of 100 speakers out of the original 250. This selection balances gender representation with 47 male and 53 female speakers and encompasses a total of 1,511 audio files, amounting to approximately 4.96 hours of speech data.
We enhance the original single-channel speech data by creating two-channel versions with gpuRIR [25], arranging two linear microphones at distances ranging from 0.05 to 0.09 meters. The gpuRIR is a free and open-source Python library for room impulse response (RIR) simulation using the image source method with GPU acceleration. The details of the simulated room environments are specified in Table 1. To ensure sound sources are well-distributed across the room, we first simulate the microphone center anywhere inside the room. Then we randomly choose a distance such that the source is inside the room to finally simulate the RIR. Based on the RIR, we generate the mixture data for each original single-channel speech data every 5 degrees from 0 to 180 degrees, which expands the audio files 37 times.
3.2 Metrics
For evaluating our models, we utilize two primary metrics: mean absolute error (MAE) and accuracy (ACC). MAE measures the average magnitude of the errors in a set of predictions, without considering their direction. On the other hand, ACC is used to determine the percentage of predictions that are exactly correct or within a certain degree of tolerance. We assess accuracy at three levels of tolerance: Within a 5-degree, 10-degree, and 15-degree error margin. These tiers of tolerance allow us to understand not only the model’s precision but also its practical effectiveness.
3.3 Incremental learning setting
In our incremental learning approach, we divide the dataset into five parts according to the varying distances between the microphones, with each partition representing a unique task. This allows us to treat the problem as a series of incremental learning tasks, each corresponding to a specific microphone spacing. We compare our approach with several baseline methods:
  • Multicondition: This involves training five separate models, each on a different subset of the data, with each model specializing in a particular microphone spacing.
  • Joint: A single model is trained on the entire dataset, encompassing all microphone spacings together. This method aims to learn a general representation that performs well across all the tasks.
  • Finetune: Starting with a model pre-trained on the initial task, we sequentially fine-tune it on each subsequent task. This approach adapts the model incrementally, but may be susceptible to forget previously learned information when moving to new tasks each time.
(a) ACC ±$\pm 5$
(b) ACC ±$\pm 10$
(c) ACC ±$\pm 15$
Figure 2: Comparative performance of various DOA estimation methods across different microphone spacings, showing ACC within ±$\pm 5$, ±$\pm 10$, and ±$\pm 15$-degree tolerances.
3.4 Implementation details
For training each task within our DOA-PNN framework, we consistently apply 100 epochs with an early stopping criterion set to activate after 20 epochs if there’s no improvement in performance. We employ the AdamW [26] optimizer, a variant of the Adam optimizer that includes weight decay for regularization, with a learning rate of 0.001. Additionally, we apply a learning rate scheduler, specifically StepLR, to adjust the learning rate during training.
For the STFT computations, we set the frame size to 32 milliseconds and the hop size to 10 milliseconds. The frequency range for the STFT is confined between a minimum of 100 Hz and a maximum of 8000 Hz, using an n_fft value of 512, which determines the length of the windowed signal after padding with zeros. This STFT configuration is designed to capture a wide range of audio frequencies suitable for accurately determining the DOA. Gaussian function constant $\sigma$ we set as 8 in our experiments.
Table 2: Performance comparison across different DOA estimation methods, highlighting MAE, ACC within various degree tolerances, and model parameters.
Methods MAE ACC(±$\pm 5$) ACC(±$\pm 10$) ACC(±$\pm 15$) Params
Finetune 17.8 33.9 48.0 58.4 0.9M
Joint 11.9 54.0 66.7 74.3 0.9M
Multicondition 7.6 78.6 84.8 87.5 4.8M
Proposed DOA-PNN 10.1 72.9 82.7 86.1 1.6M
4 Results and Discussions
Delving into the data for DOA estimation across varying microphone spacings as Figure 2, we can observe how each method adapts to the acoustic challenge. The performance of finetune method degrades as microphone spacing increases, indicating a potential forgetting across previously learned configurations. The rationale behind this is attributed due to strong focus on the initial training data, making it less flexible when conditions change. On looking to the perfroamnce for joint training, we find that it is more resilience compared to finetune, yet it does not consistently maintain high performance as spacing widens. It is possible that the single-model approach is not able to capture the nuances needed for precise localization at greater spacings. The multicondition method demonstrates a more robust performance across different spacings, likely benefiting from its strategy of training distinct models for each spacing scenario. This suggests that dedicated learning per spacing scenario captures specific acoustic features that general models might miss.
Our proposed DOA-PNN shows a promising ability to sustain higher accuracy levels, particularly within the most challenging ±$\pm 5$-degree margin. It implies that our progressive learning strategy, which builds upon knowledge from previous configurations, is effective in developing an understanding that transfers well to new spacing scenarios. The proposed DOA-PNN offers a balance between accuracy and model complexity, suggesting it to be a versatile and efficient choice for real-world applications where microphone spacing can vary.
On examining the comparative results of different methods for DOA estimation as given in Table 2, we find each method has its strengths and limitations. The finetune approach, while the most straightforward in terms of adapting pre-existing models to new tasks, seems to lag in precision. It presents the highest MAE, suggesting that it might not preserve the nuances of DOA estimation in different microphone spacing as effectively as the other methods. The joint training method improves upon finetune significantly, indicating that learning from a collective dataset of all tasks simultaneously offers a better understanding of the DOA estimation problem. However, this method still does not reach the top performance in accuracy, especially in ±$\pm 5$-degree range. Multicondition, which employs a separate model for each condition, stands out in accuracy. It’s clear that specializing in individual conditions allows for a finer understanding of each specific DOA estimation task, but this comes at the cost of a larger overall model size, as reflected by the parameter count.
Our Proposed DOA-PNN method maintains a balance between precision and efficiency. It achieves comparable accuracy, nearly matching the best-performing method in the ±$\pm 15$-degree range and surpassing it within the ±$\pm 10$-degree margin. It achieves this while maintaining a moderate model size, suggesting an efficient use of parameters. This indicates that our progressive approach, which incorporates learning from successive tasks, effectively consolidates knowledge without increasing complexity.
We foresee several promising avenues for future work. While our current study focused on two-microphone linear arrays to estimate the azimuth angle, we plan to expand our research to encompass more complex acoustic configurations and to address elevation angle estimation for a full spatial representation [27]. Moreover, integrating noise robustness factors into DOA estimation is a critical next step, acknowledging that real-world applications often operate in acoustically challenging environments [28]. These enhancements will aim to bolster the practicality of DOA-PNN, providing comprehensive solutions for 3D sound localization in noisy, real-life settings.
5 Conclusions
In this study, we presented DOA-PNN, an innovative approach that advances the direction of arrival estimation within the continual learning paradigm. DOA-PNN synergizes task-specific network instantiation with progressive learning principles, addressing the diversity in microphone configurations with a novel scaling mechanism and lateral connections for knowledge retention. Demonstrated across simulated acoustic settings, DOA-PNN consistently achieved high accuracy, underscoring its capacity for real-world applicability in dynamic auditory scenes. The model’s efficiency and adaptability suggest its potential as a foundational framework for future research in sound localization and acoustic signal processing, encouraging the development of sophisticated, scalable audio analysis technologies.

### v2

@@
 The studies reveal its capability to maintain performance with minimal parameter increase, presenting an efficient solution for DOA estimation.
-Index Terms—  DOA Estimation, Sound Source Localization, Incremental Learning, Configurable Model
-1 Introduction
-Sound source localization (SSL) [1] aims to identify the origins of sound sources automatically by analyzing signals captured through an array of microphones.
+Index Terms:
+DOA Estimation, Sound Source Localization, Incremental Learning, Configurable Model
+I Introduction
+Sound source localization (SSL) [1] aims to identify the origins of sound sources by analyzing signals captured through an array of microphones.
 At the cornerstone of SSL lies the direction of arrival (DOA) estimation, a process crucial for calculating the angles at which sounds reach the microphone array.
-This spatial information is necessary for a broad range of speech-related applications, including but not limited to, speech separation [2], speaker extraction [3], speech recognition [4], and speech enhancement [5].
+This spatial information is necessary for speech-related applications [2, 3, 4, 5].
 These applications rely on the precise localization of sound sources to significantly improve their performance and functionality.
 Consequently, the development of more sophisticated DOA estimation techniques has become a key focus within the research community to achieve greater robustness and adaptability in real-world settings.
-Classical signal processing techniques like generalized cross-correlation (GCC) [6], multiple signal classification (MUSIC) [7], and steered response power (SRP) based methods, particularly SRP-PHAT [8], which is obtained from the SRP by applying phase transform (PHAT) whitening, have been foundational due to their simplicity and versatility.
+Classical signal processing techniques like generalized cross-correlation (GCC) [6], multiple signal classification (MUSIC) [7], and steered response power (SRP) based methods, particularly SRP-PHAT [8], which is obtained from the SRP by applying phase transform (PHAT) whitening.
 They are available to adapt to different microphone array configurations.
-However, their effectiveness decreases in noisy and reverberant environments, which leads to exploration of deep learning solutions.
+However, their effectiveness decreases in noisy and reverberant environments compared with deep learning solutions.
 A significant milestone in this exploration is the adoption of deep learning architectures such as convolutional neural networks (CNN) [9, 10, 11, 12, 13].
 However, deep learning models used for DOA estimation face a unique challenge: the misalignment between training and testing environments, especially concerning the microphone array’s configuration.
-Such mismatches, particularly changes in microphone inter-distances, precipitate a decline in model efficacy, necessitating re-training for new configurations.
-Re-training is not always feasible and requires extra time or computation resources.
-This highlights an essential area for ongoing research and innovation of deep learning-based DOA estimation, to enhance model flexibility and accuracy across diverse acoustic settings.
-Recent advancements in deep DOA estimation [14] tackle array configuration mismatching by framing DOA as a classification problem within each time-frequency (TF) bin, using high-resolution images from narrow-band SRP-PHAT for input.
-With such high-level localization cues, relatively simple neural networks can be used.
-Although these methods adapt to the unseen array setups relying on specific spatial features, they do not address the issue when incrementally adjusting to new acoustic configurations.
+Such mismatches precipitate a decline in model efficacy, necessitating re-training for new configurations.
+Re-training always requires extra time and computation resources.
+This highlights an essential area for ongoing deep learning-based DOA estimation research, to enhance model accuracy across diverse acoustic settings.
+Although some methods [14, 15] adapt to the variable array setups relying on specific spatial features, they do not address more complex acoustic configurations like the difference absorption and reverberation time.
 Along the diverse configurations, one fixed method faces forgetting because the new knowledge overrides the existing knowledge.
-Various research recently incorporated continual learning (CL) [15] to continuously learn new knowledge while retaining and reusing previously learned knowledge in keyword spotting [16], audio classification [17], and sound event detection.
+Various research recently incorporated continual learning (CL) [16] to continuously learn new knowledge while retaining previously learned knowledge in audio classification [17, 18].
 Task Incremental Learning (TIL)  [19] is a subset of CL methods.
 The TIL approach enables the model to perform new tasks efficiently with the same number of classes in each task without losing prior performance.
-Inspired by [19, 20], we proposed to use TIL to gradually learn new information for DOA estimation using simple, widely used features like the short-time Fourier transform (STFT), all while keeping the valuable knowledge it has already gained from the existing acoustic environments and configurations.
+Inspired by [20, 21, 22], we proposed to use TIL to gradually learn new information for DOA estimation using simple, widely used features like the short-time Fourier transform (STFT), all while keeping the valuable knowledge it has already gained from the existing acoustic environments and configurations.
 In this research, we introduce DOA-PNN, a novel progressive neural network based on the TIL method for DOA estimation, crafted to solve the issue of adaptation from source to target configurations.
@@
 Contrary to traditional finetuning methods, DOA-PNN utilizes a collection of pre-trained sub-networks, establishing lateral connections that extract the crucial features for new tasks, thereby enriching feature presentation.
-To the best of our knowledge, we are the first to apply continual learning techniques to adapt DOA estimation to various acoustic configurations.
+To the best of our knowledge, this is the first work to apply continual learning techniques to adapt DOA estimation to various acoustic configurations.
 Demonstrated through extensive testing on a simulated dataset for DOA estimation, DOA-PNN exhibits competitive performance on unfamiliar configurations with less parameter increase and no reliance on previous data buffers.
 Figure 1: (a) Architecture of SSNet (b) Proposed DOA-PNN framework.
-2 Proposed Method
-2.1 Continual learning for DOA estimation
-In this work, applying continual learning to DOA estimation involves receiving a series of $T$ distinct DOA estimation tasks (under different acoustic configurations) in sequence and optimizing the performance across all tasks with less negative impact of catastrophic forgetting.
+II Proposed Method
+II-A Continual learning for DOA estimation
+In this work, applying continual learning to DOA estimation involves receiving a series of $T$ distinct tasks (under different acoustic configurations) in sequence and optimizing the performance across all tasks with less negative impact of catastrophic forgetting.
 For each task $\tau_{t}$, with $t\leq T$, we handle a set of training data consisting of pairs $(x_{t},y_{t})$, where $x_{t}$ represents the acoustic signals and $y_{t}$ are the corresponding DOA labels, sampled from a distribution $D^{T}$.
@@
 where $L_{doa}$ represents a classification loss function, such as mean squared error (MSE) loss, and $F_{t}(x_{t};\mu_{t})$ is the DOA estimation model parameterized by $\mu_{t}$.
-During the training phase, we aim to find the optimal parameters $\mu_{t}^{\prime}$ that perform well across all $T$ tasks.
+During the training, we aim to find the optimal parameters $\mu_{t}^{\prime}$ that perform well across all $T$ tasks.
 This is challenging since parameters learned from previous task $\tau_{t}$ are prone to be overwritten when learning the subsequent task $\tau_{t+1}$, leading to catastrophic forgetting.
-2.2 Network backbone for DOA estimation
+II-B Network backbone for DOA estimation
 We adopt a spatial spectrum prediction module using the spatial spectrum network (SSNet) [23, 24, 25] for DOA estimation.
@@
 This $v$ feature is then used as the input to SSNet for predicting the spatial spectrum $\hat{p}$.
-The network structure of SSNet is shown in Figure 1 (a) that depicts SSNet includes layers to simplify and analyze sound features.
+The network structure of SSNet is shown in Fig.
+1 (a) that depicts SSNet includes layers to simplify and analyze sound features.
 It uses two convolutional layers to reduce frequency details and five residual blocks to pick up intricate sound features.
@@
 We then train SSNet by computing the MSE loss between the segment-wise output spatial spectrum from the model and the ground truth.
-2.3 Proposed DOA-PNN
-Drawing on the concept of progressive neural networks, we introduce a novel progressive continual learning approach specifically designed for deep DOA estimation, depicted in Figure 1(b).
+II-C Proposed DOA-PNN
+Drawing on the concept of progressive neural networks, we introduce a novel progressive continual learning approach specifically designed for deep DOA estimation, depicted in Fig.
+1(b).
 Our method, termed DOA-PNN, is composed of two principal elements: a task-specific sub-network creator for individual tasks and a lateral connection set bridging a suite of sub-networks.
@@
 To integrate these multiple outputs effectively, each sub-network is equipped with an adapter—a 1$\times$1 convolutional layer—tailored to its current context.
-This adapter transforms the combined features into the final output, thereby enriching the feature used for classification with the accumulated knowledge from the predecessors.
+This 1$\times$1 convolutional adapter transforms the combined features into the final output, thereby enriching the feature used for classification with the accumulated knowledge from the predecessors.
 A gap-aware layer scaling mechanism adjusts the complexity of the SSNet feature extractor’s residual blocks based on a consistent parameter reflecting the minimum angle error tolerance.
-For instance, if the smallest acceptable angle error is 5 degrees, we streamline the residual blocks from five down to one.
-This scaling ensures the network remains efficient and tailored to the specific precision requirements of the DOA estimation task.
+If the smallest acceptable error is 5 degrees, we streamline the residual blocks from five down to one.
+This manual scaling ensures the network remains efficient and tailored to the precision requirements of the DOA estimation.
 In this architecture, for each new learning task labeled as $\tau_{t}$, we compute the real and imaginary parts of the complex STFT of audio segments to serve as input data.
@@
 What sets the $t^{{\text{th}}}$ sub-network apart is its ability to access the shared memory of its predecessors through lateral connections, integrating all previous knowledge.
-This structure not only facilitates effective knowledge transfer across various tasks, but also promotes faster convergence during the training process.
 When it comes to inference application, the DOA-PNN model selects the appropriate $t^{{\text{th}}}$ sub-network based on the specific task at hand for accurate and reliable evaluation.
 This approach ensures that our model remains adaptable and quick to respond in diverse acoustic settings, showcasing the potential of progressive continual learning strategy in the real-world environments of DOA estimation.
-3 Experiments
-Table 1: The parameter settings of simulated audio
+III Experiments
+TABLE I: Parameter settings of simulated audio
 Room Size [L, W, H] Absorption Tmax Num_images
 [10,8,5], [5,4,3], [4,2,2] 0.2-0.8 0.2-0.6 [10,10,10]
-3.1 Dataset
+III-A Dataset
 We use the LibriSpeech [26] train-clean-100 subset, which comprises high-fidelity, single-channel speech recordings sampled at 16 kHz.
@@
 The gpuRIR is a free and open-source Python library for room impulse response (RIR) simulation using the image source method with GPU acceleration.
-The details of the simulated room environments are specified in Table 1.
+The details of the simulated room environments are specified in Table I.
 To ensure sound sources are well-distributed across the room, we first simulate the microphone center anywhere inside the room.
@@
 Based on the RIR, we generate the mixture data for each original single-channel speech data every 5 degrees from 0 to 180 degrees, which expands the audio files 37 times.
-3.2 Metrics
+III-B Metrics
 For evaluating our models, we utilize two primary metrics: mean absolute error (MAE) and accuracy (ACC).
@@
 These tiers of tolerance allow us to understand not only the model’s precision but also its practical effectiveness.
-3.3 Incremental learning setting
-In our incremental learning approach, we divide the dataset into five parts according to the varying distances between the microphones, with each partition representing a unique task.
-This allows us to treat the problem as a series of incremental learning tasks, each corresponding to a specific microphone spacing.
-We compare our approach with several baseline methods:
-  • Multicondition: This involves training five separate models, each on a different subset of the data, with each model specializing in a particular microphone spacing.
-  • Joint: A single model is trained on the entire dataset, encompassing all microphone spacings together. This method aims to learn a general representation that performs well across all the tasks.
-  • Finetune: Starting with a model pre-trained on the initial task, we sequentially fine-tune it on each subsequent task.
-This approach adapts the model incrementally, but may be susceptible to forget previously learned information when moving to new tasks each time.
 (a) ACC ±$\pm 5$
@@
 (c) ACC ±$\pm 15$
-Figure 2: Comparative performance of various DOA estimation methods across different microphone spacings, showing ACC within ±$\pm 5$, ±$\pm 10$, and ±$\pm 15$-degree tolerances.
-3.4 Implementation details
+Figure 2: Comparative inference performance of various DOA estimation methods across different microphone spacings, showing ACC within three tolerances.
+III-C Incremental learning setting
+In our incremental learning approach, we divide the dataset into five parts according to the varying distances between the microphones, with each partition representing a unique task.
+This allows us to treat the problem as a series of incremental learning tasks, each corresponding to a microphone spacing.
+We compare our approach with several baseline methods:
+  • Finetune: Starting with a model pre-trained on the initial task, we sequentially fine-tune it on each subsequent task.
+This approach adapts the model incrementally, but may be susceptible to forget previously learned information when moving to new tasks each time.
+  • Joint: A single model is trained on the entire dataset, encompassing all microphone spacings together. This method aims to learn a general representation that performs well across all the tasks.
+  • Ensemble: This involves training five separate models, each on a different subset of the data, with each model specializing in a particular microphone spacing.
+III-D Implementation details
 For training each task within our DOA-PNN framework, we consistently apply 100 epochs with an early stopping criterion set to activate after 20 epochs if there’s no improvement in performance.
@@
 Gaussian function constant $\sigma$ we set as 8 in our experiments.
-Table 2: Performance comparison across different DOA estimation methods, highlighting MAE, ACC within various degree tolerances, and model parameters.
+TABLE II: Performance comparison across different DOA estimation methods, highlighting MAE, ACC within various degree tolerances, and model parameters.
 Methods MAE ACC(±$\pm 5$) ACC(±$\pm 10$) ACC(±$\pm 15$) Params
@@
 Joint 11.9 54.0 66.7 74.3 0.9M
-Multicondition 7.6 78.6 84.8 87.5 4.8M
+Ensemble 7.6 78.6 84.8 87.5 4.8M
 Proposed DOA-PNN 10.1 72.9 82.7 86.1 1.6M
-4 Results and Discussions
-Delving into the data for DOA estimation across varying microphone spacings as Figure 2, we can observe how each method adapts to the acoustic challenge.
-The performance of finetune method degrades as microphone spacing increases, indicating a potential forgetting across previously learned configurations.
-The rationale behind this is attributed due to strong focus on the initial training data, making it less flexible when conditions change.
-On looking to the perfroamnce for joint training, we find that it is more resilience compared to finetune, yet it does not consistently maintain high performance as spacing widens.
-It is possible that the single-model approach is not able to capture the nuances needed for precise localization at greater spacings.
-The multicondition method demonstrates a more robust performance across different spacings, likely benefiting from its strategy of training distinct models for each spacing scenario.
-This suggests that dedicated learning per spacing scenario captures specific acoustic features that general models might miss.
+IV Results and Discussions
+We test the methods for DOA estimation across varying microphone spacings as Fig.
+2, we can observe how each method adapts to the acoustic challenge.
+The finetune method incrementally trained from 5cm to 9cm only performs well for the recent task 9 cm indicating a forgetting across previously learned configurations.
+On looking to the performance for joint training, we find that it is more resilience compared to finetune, yet it does not maintain high performance as spacing widens.
+It is possible that the single-model approach is not able to capture the nuances needed for localization at variable spacings.
+The ensemble method demonstrates a robust performance across different spacings as the upper bound.
+It benefits from training distinct models for each spacing scenario.
 Our proposed DOA-PNN shows a promising ability to sustain higher accuracy levels, particularly within the most challenging ±$\pm 5$-degree margin.
-It implies that our progressive learning strategy, which builds upon knowledge from previous configurations, is effective in developing an understanding that transfers well to new spacing scenarios.
-The proposed DOA-PNN offers a balance between accuracy and model complexity, suggesting it to be a versatile and efficient choice for real-world applications where microphone spacing can vary.
-On examining the comparative results of different methods for DOA estimation as given in Table 2, we find each method has its strengths and limitations.
+It implies that our progressive learning strategy, which builds upon knowledge from previous configurations transfers well to new spacing scenarios.
+The proposed DOA-PNN offers a balance between accuracy and model complexity, suggesting it to be an efficient choice for real-world applications where microphone spacing can vary.
+On examining the comparative results of different methods for DOA estimation as given in Table II, we find each method has its strengths and limitations.
 The finetune approach, while the most straightforward in terms of adapting pre-existing models to new tasks, seems to lag in precision.
 It presents the highest MAE, suggesting that it might not preserve the nuances of DOA estimation in different microphone spacing as effectively as the other methods.
-The joint training method improves upon finetune significantly, indicating that learning from a collective dataset of all tasks simultaneously offers a better understanding of the DOA estimation problem.
+The joint training method improves upon finetune significantly, indicating that learning from a collective dataset of all tasks offers a better understanding of the DOA estimation problem.
 However, this method still does not reach the top performance in accuracy, especially in ±$\pm 5$-degree range.
-Multicondition, which employs a separate model for each condition, stands out in accuracy.
+Ensemble, which employs a separate model for each condition, stands out in accuracy.
 It’s clear that specializing in individual conditions allows for a finer understanding of each specific DOA estimation task, but this comes at the cost of a larger overall model size, as reflected by the parameter count.
 Our Proposed DOA-PNN method maintains a balance between precision and efficiency.
-It achieves comparable accuracy, nearly matching the best-performing method in the ±$\pm 15$-degree range and surpassing it within the ±$\pm 10$-degree margin.
-It achieves this while maintaining a moderate model size, suggesting an efficient use of parameters.
+It achieves comparable accuracy, nearly matching the best-performing method in the ±$\pm 15$-degree range and outperforming it within the ±$\pm 10$-degree margin.
+It achieves this while maintaining a moderate model size.
 This indicates that our progressive approach, which incorporates learning from successive tasks, effectively consolidates knowledge without increasing complexity.
-We foresee several promising avenues for future work.
-While our current study focused on two-microphone linear arrays to estimate the azimuth angle, we plan to expand our research to encompass more complex acoustic configurations and to address elevation angle estimation for a full spatial representation [27].
+We foresee several avenues for future work.
+While we currently focus on two-microphone linear arrays to estimate the azimuth angle, we plan to expand our research to encompass more complex acoustic configurations and to address elevation angle estimation for a full spatial representation [29].
 Moreover, integrating noise robustness factors into DOA estimation is a critical next step, acknowledging that real-world applications often operate in acoustically challenging environments [30].
 These enhancements will aim to bolster the practicality of DOA-PNN, providing comprehensive solutions for 3D sound localization in noisy, real-life settings.
-5 Conclusions
-In this study, we presented DOA-PNN, an innovative approach that advances the direction of arrival estimation within the continual learning paradigm.
+V Conclusions
+In this study, we presented DOA-PNN, an approach that advances the direction of arrival estimation within continual learning.
 DOA-PNN synergizes task-specific network instantiation with progressive learning principles, addressing the diversity in microphone configurations with a novel scaling mechanism and lateral connections for knowledge retention.

### v3

@@
-Configurable DOA Estimation using Incremental Learning
+Where’s That Voice Coming? Continual Learning for Sound Source Localization
 Abstract
-This study introduces a progressive neural network (PNN) model for direction of arrival (DOA) estimation, DOA-PNN, addressing the challenge due to catastrophic forgetting in adapting dynamic acoustic environments.
-While traditional methods such as GCC, MUSIC, and SRP-PHAT are effective in static settings, they perform worse in noisy, reverberant conditions.
-Deep learning models, particularly CNNs, offer improvements but struggle with a mismatch configuration between the training and inference phases.
-The proposed DOA-PNN overcomes these limitations by incorporating task incremental learning of continual learning, allowing for adaptation across varying acoustic scenarios with less forgetting of previously learned knowledge.
-Featuring task-specific sub-networks and a scaling mechanism, DOA-PNN efficiently manages parameter growth, ensuring high performance across incremental microphone configurations.
-We study DOA-PNN on a simulated data under various mic distance based microphone settings.
-The studies reveal its capability to maintain performance with minimal parameter increase, presenting an efficient solution for DOA estimation.
+Sound source localization (SSL) is essential for many speech-processing applications.
+Deep learning models have achieved high performance, but often fail when the training and inference environments differ.
+Adapting SSL models to dynamic acoustic conditions faces a major challenge: catastrophic forgetting.
+In this work, we propose an exemplar-free continual learning strategy for SSL (CL-SSL) to address such a forgetting phenomenon.
+CL-SSL applies task-specific sub-networks to adapt across diverse acoustic environments while retaining previously learned knowledge.
+It also uses a scaling mechanism to limit parameter growth, ensuring consistent performance across incremental tasks.
+We evaluated CL-SSL on simulated data with varying microphone distances and real-world data with different noise levels.
+The results demonstrate CL-SSL’s ability to maintain high accuracy with minimal parameter increase, offering an efficient solution for SSL applications.
 Index Terms:
-DOA Estimation, Sound Source Localization, Incremental Learning, Configurable Model
+DOA Estimation, Sound Source Localization, Incremental Learning, Continual Learning
 I Introduction
-Sound source localization (SSL) [1] aims to identify the origins of sound sources by analyzing signals captured through an array of microphones.
+Sound source localization (SSL) [1] aims to identify the origins of speech sources by analyzing signals captured through an array of microphones.
 At the cornerstone of SSL lies the direction of arrival (DOA) estimation, a process crucial for calculating the angles at which sounds reach the microphone array.
-This spatial information is necessary for speech-related applications [2, 3, 4, 5].
-These applications rely on the precise localization of sound sources to significantly improve their performance and functionality.
-Consequently, the development of more sophisticated DOA estimation techniques has become a key focus within the research community to achieve greater robustness and adaptability in real-world settings.
+This spatial information is necessary for speech-related applications [2, 3, 4, 5] such as speech enhancement and far-field automatic speech recognition.
+These applications rely on the precise localization of speech sources to significantly improve their performance and functionality.
+Consequently, the development of more sophisticated SSL techniques has become a key focus within the research community to achieve greater robustness and adaptability in real-world settings.
 Classical signal processing techniques like generalized cross-correlation (GCC) [6], multiple signal classification (MUSIC) [7], and steered response power (SRP) based methods, particularly SRP-PHAT [8], which is obtained from the SRP by applying phase transform (PHAT) whitening.
@@
 A significant milestone in this exploration is the adoption of deep learning architectures such as convolutional neural networks (CNN) [9, 10, 11, 12, 13].
-However, deep learning models used for DOA estimation face a unique challenge: the misalignment between training and testing environments, especially concerning the microphone array’s configuration.
+However, deep learning models used for SSL face a unique challenge: the misalignment between training and testing environments, especially concerning the microphone array’s configuration and acoustic environments.
 Such mismatches precipitate a decline in model efficacy, necessitating re-training for new configurations.
 Re-training always requires extra time and computation resources.
-This highlights an essential area for ongoing deep learning-based DOA estimation research, to enhance model accuracy across diverse acoustic settings.
+This highlights an essential area for ongoing deep learning-based SSL research, to enhance model accuracy across diverse acoustic settings.
 Although some methods [14, 15] adapt to the variable array setups relying on specific spatial features, they do not address more complex acoustic configurations like the difference absorption and reverberation time.
 Along the diverse configurations, one fixed method faces forgetting because the new knowledge overrides the existing knowledge.
-Various research recently incorporated continual learning (CL) [16] to continuously learn new knowledge while retaining previously learned knowledge in audio classification [17, 18].
-Task Incremental Learning (TIL)  [19] is a subset of CL methods.
-The TIL approach enables the model to perform new tasks efficiently with the same number of classes in each task without losing prior performance.
-Inspired by [20, 21, 22], we proposed to use TIL to gradually learn new information for DOA estimation using simple, widely used features like the short-time Fourier transform (STFT), all while keeping the valuable knowledge it has already gained from the existing acoustic environments and configurations.
-In this research, we introduce DOA-PNN, a novel progressive neural network based on the TIL method for DOA estimation, crafted to solve the issue of adaptation from source to target configurations.
-The core of DOA-PNN’s architecture is task-specific sub-networks , each engineered to memorize knowledge from previous configurations, enabling smooth transitions across diverse acoustic environments while maintaining accuracy on previous tasks.
+This problem is known as catastrophic forgetting.
+Recently, continual learning (CL) [16] has received notable attention for solving catastrophic forgetting.
+Various works recently incorporated CL to continuously learn new knowledge while retaining previously learned knowledge in speech processing [17, 18, 19].
+We can classify existing CL methods into two categories based on whether to use a replay buffer, which stores extra data: exemplar-free methods and replay-based methods.
+Exemplar-free methods protect parameters learned from previous tasks through loss regularization or dynamic modules without access to any historical data in existing tasks.
+Replay-based methods store historical data as replay exemplars in a memory buffer.
+Since SSL applications are typically deployed on edge devices with limited memory, we propose using exemplar-free methods [20, 21, 22].
+These methods allow SSL to gradually learn new information using simple, widely adopted features like the short-time Fourier transform (STFT) while retaining valuable knowledge from existing acoustic environments and configurations.
+In this research, we introduce CL-SSL, a novel exemplar-free continual learning method for SSL, crafted to solve the issue of adaptation from source to target configurations and environments.
+The core of CL-SSL’s architecture is task-specific sub-networks , each engineered to memorize knowledge from previous configurations or environments, enabling smooth transitions across diverse acoustic environments while maintaining accuracy on previous tasks.
 A scaling mechanism is integrated into these sub-networks to limit parameter growth by reducing the extra layers.
-Contrary to traditional finetuning methods, DOA-PNN utilizes a collection of pre-trained sub-networks, establishing lateral connections that extract the crucial features for new tasks, thereby enriching feature presentation.
-To the best of our knowledge, this is the first work to apply continual learning techniques to adapt DOA estimation to various acoustic configurations.
-Demonstrated through extensive testing on a simulated dataset for DOA estimation, DOA-PNN exhibits competitive performance on unfamiliar configurations with less parameter increase and no reliance on previous data buffers.
-Figure 1: (a) Architecture of SSNet (b) Proposed DOA-PNN framework.
-II Proposed Method
-II-A Continual learning for DOA estimation
-In this work, applying continual learning to DOA estimation involves receiving a series of $T$ distinct tasks (under different acoustic configurations) in sequence and optimizing the performance across all tasks with less negative impact of catastrophic forgetting.
-For each task $\tau_{t}$, with $t\leq T$, we handle a set of training data consisting of pairs $(x_{t},y_{t})$, where $x_{t}$ represents the acoustic signals and $y_{t}$ are the corresponding DOA labels, sampled from a distribution $D^{T}$.
-Our goal is to minimize the cumulative loss $L_{tot}$, which aggregates the sub-losses $L_{doa}$ for all $T$ tasks, formulated as:
-$L_{tot}=\sum_{t=0}^{T}E_{(x_{t},y_{t})\sim D^{T}}[L_{doa}(F_{t}(x_{t};\mu_{t}),y_{t})]$ (1)
-where $L_{doa}$ represents a classification loss function, such as mean squared error (MSE) loss, and $F_{t}(x_{t};\mu_{t})$ is the DOA estimation model parameterized by $\mu_{t}$.
+Contrary to traditional finetuning methods, CL-SSL utilizes a collection of pre-trained sub-networks, establishing lateral connections that extract the crucial features for new tasks, thereby enriching feature presentation.
+To the best of our knowledge, this is the first work to apply continual learning techniques to adapt SSL to various acoustic configurations.
+Demonstrated through extensive testing on the simulated and real-world data, CL-SSL exhibits competitive performance on unfamiliar configurations and acoustic environments with less parameter increase and no reliance on previous data buffers.
+Figure 1: (a) Architecture of SSNet (b) Proposed CL-SSL framework.
+II Related Works
+Recently, several studies have explored the application of continual learning in speech-processing tasks.
+In [23], the authors proposed a hyper-gradient-based exemplar strategy for task-oriented dialogue systems, which selects key exemplars for periodic retraining.
+Xiao et al.
+[22] developed an independent unsupervised learning framework with a distillation loss to incorporate new sound classes while maintaining the consistency of sound event detection across incremental tasks.
+The authors of [24] introduced incremental spoken keyword spotting, using model predictions to distill past experiences throughout training.
+However, these methods do not address the issue of catastrophic forgetting in SSL tasks.
+In addition, most approaches rely on exemplars from existing tasks, making them less suitable for on-device applications with memory constraints.
+In a CL-based SSL task reported in [25], the authors treat azimuth angles as discrete classes, which may not reflect in real-world scenarios as splitting the range $360^{\circ}$ into multiple tasks creates artificial constraints and reduces practical applicability.
+III Proposed Method
+III-A Continual learning for sound source localization
+In this work, applying continual learning to SSL involves receiving a series of $T$ distinct tasks (under different configurations and environments) in sequence and optimizing the performance across all tasks with less catastrophic forgetting.
+For each task $\tau_{t}$, with $t\leq T$, we handle a set of training data consisting of pairs $(x_{t},y_{t})$, where $x_{t}$ represents the acoustic signals and $y_{t}$ are the corresponding direction labels, sampled from a distribution $D^{T}$.
+Our goal is to minimize the cumulative loss $L_{tot}$, which aggregates the sub-losses $L_{ssl}$ for all $T$ tasks, formulated as:
+$L_{tot}=\sum_{t=0}^{T}E_{(x_{t},y_{t})\sim D^{T}}[L_{ssl}(F_{t}(x_{t};\mu_{t}),y_{t})]$ (1)
+where $L_{ssl}$ represents a classification loss function, such as mean squared error (MSE) loss, and $F_{t}(x_{t};\mu_{t})$ is the SSL model parameterized by $\mu_{t}$.
 During the training, we aim to find the optimal parameters $\mu_{t}^{\prime}$ that perform well across all $T$ tasks.
 This is challenging since parameters learned from previous task $\tau_{t}$ are prone to be overwritten when learning the subsequent task $\tau_{t+1}$, leading to catastrophic forgetting.
-II-B Network backbone for DOA estimation
-We adopt a spatial spectrum prediction module using the spatial spectrum network (SSNet) [23, 24, 25] for DOA estimation.
-We consider 2 microphone linear arrays as an instance in this section, although our proposed method can be employed for any number of microphone positions.
-This system learns complex relationships from audio data to predict, where the sound could be coming from at every degree in a half-circle of the linear microphone array.
-To perform this, we combine sound segments from different microphones and transform them into complex STFT.
+III-B Network backbone for sound source localization
+We adopt a spatial spectrum prediction module using the spatial spectrum network (SSNet) [26, 27, 28] for SSL.
+We consider two microphone linear arrays as an instance in this section, although our proposed method can be employed for any number of microphone positions.
+This system learns complex relationships from audio data for predictions, where the sound could be coming from at every degree in a half-circle of the linear microphone array.
+To perform this, we combine sound segments from two microphones and transform them into complex STFT.
 We consider the real as well as the imaginary part of the complex STFT instead of the phase and power and then combine this data to form a feature representation $v$.
@@
 The network structure of SSNet is shown in Fig.
-1 (a) that depicts SSNet includes layers to simplify and analyze sound features.
+1 (a) which depicts SSNet includes layers to simplify and analyze sound features.
 It uses two convolutional layers to reduce frequency details and five residual blocks to pick up intricate sound features.
-Then, it has a layer that focuses on the DOA estimation and rearranges the sound information in preparation for the final step.
-The result spatial spectrum is a 181-point map.
+Then, it has a layer that focuses on the SSL and rearranges the sound information in preparation for the final step.
+The result spatial spectrum is an 181-point map.
 Following the work in [28], we use a probability-based method to determine the likelihood of a sound coming from any of the 180 directions.
-This part of SSNet assigns a probability to each direction based on how closely it matches the ground truth DOA.
+This part of SSNet assigns a probability to each direction based on how closely it matches the ground truth direction.
 Specifically, each element of the encoded $181$-dimensional vector $p(\theta_{i})$ is assigned to a particular azimuth direction $\theta_{i}\in\{0^{\circ},1^{\circ},\dots,180^{\circ}\}$.
-The values of the vector follow a Gaussian distribution that maximizes at the ground truth DOA, which is defined as follows:
+The values of the vector follow a Gaussian distribution that maximizes at the ground truth direction, which is defined as follows:
 $p(\theta_{i})=\begin{cases}\exp\left(-\frac{d(\theta_{i},\theta_{y})^{2}}{2\sigma^{2}}\right),&\text{if }|\Theta|=0\\ 1,&\text{if }|\Theta|=1\end{cases}$ (2)
-where $|\Theta|$ is the number of sources, $\theta^{\prime}$ is the ground truth DOA of one source, $\sigma$ is a predefined constant that controls the width of the Gaussian function and $d(\dots,\dots)$ denotes the angular distance.
+where $|\Theta|$ is the number of sources, $\theta^{\prime}$ is the ground truth direction of one source, $\sigma$ is a predefined constant that controls the width of the Gaussian function and $d(\dots,\dots)$ denotes the angular distance.
 We then train SSNet by computing the MSE loss between the segment-wise output spatial spectrum from the model and the ground truth.
-II-C Proposed DOA-PNN
-Drawing on the concept of progressive neural networks, we introduce a novel progressive continual learning approach specifically designed for deep DOA estimation, depicted in Fig.
+In this work, we only consider the single-source localization.
+III-C Proposed CL-SSL approach
+Inspired by the progressive neural networks [20], we present a novel exemplar-free continual learning approach tailored for deep SSL, depicted in Fig.
 1(b).
-Our method, termed DOA-PNN, is composed of two principal elements: a task-specific sub-network creator for individual tasks and a lateral connection set bridging a suite of sub-networks.
-To manage the model size, we’ve introduced a gap-aware layer scaling mechanism in DOA-PNN, ensuring the parameter count remains controlled.
+Our method, termed CL-SSL, is composed of two principal elements: a task-specific sub-network creator for individual tasks and a lateral connection set bridging a suite of sub-networks.
+To manage the model size, we’ve introduced a gap-aware layer scaling mechanism in CL-SSL, ensuring the parameter count remains controlled.
 A task-specific sub-network creator is designed to generate the $t^{{\text{th}}}$ sub-network for each new task $\tau_{t}$.
-When a new acoustic configuration DOA estimation $\tau_{t}$ comes, the creator adds a sub-network with random initialization which is only the feature extractor part of SSNet.
+When a new acoustic configuration SSL $\tau_{t}$ comes, the creator adds a sub-network with random initialization which is only the feature extractor part of SSNet.
 All sub-networks share the common memory which is the classification head of SSNet for a fixed class number 181.
@@
 When the $t^{{\text{th}}}$ sub-network processes input data, it not only forwards its output to the classification layer but also receives output features from all previously established sub-networks.
-To integrate these multiple outputs effectively, each sub-network is equipped with an adapter—a 1$\times$1 convolutional layer—tailored to its current context.
+To integrate these multiple outputs effectively, each sub-network is equipped with an adapter—1$\times$1 convolutional layer—tailored to its current context.
 This 1$\times$1 convolutional adapter transforms the combined features into the final output, thereby enriching the feature used for classification with the accumulated knowledge from the predecessors.
-A gap-aware layer scaling mechanism adjusts the complexity of the SSNet feature extractor’s residual blocks based on a consistent parameter reflecting the minimum angle error tolerance.
-If the smallest acceptable error is 5 degrees, we streamline the residual blocks from five down to one.
-This manual scaling ensures the network remains efficient and tailored to the precision requirements of the DOA estimation.
+A gap-aware layer scaling mechanism adjusts the complexity of the SSNet feature extractor’s residual blocks.
+Following [21], we scale the blocks based on the minimum angle error tolerance.
+For the smallest error tolerance of 5 degrees, we reduce the residual blocks from five to one because leveraging lateral connections allows sub-networks to share capacity.
+This manual scaling constrains the growth of the model parameters for more efficient SSL applications.
 In this architecture, for each new learning task labeled as $\tau_{t}$, we compute the real and imaginary parts of the complex STFT of audio segments to serve as input data.
-Concurrently, the DOA-PNN model preserves the integrity of the previously trained $t-1$ sub-networks by freezing their parameters, preventing interference with newly acquired knowledge.
+Concurrently, the CL-SSL approach preserves the integrity of the previously trained $t-1$ sub-networks by freezing their parameters, preventing interference with newly acquired knowledge.
 The creator then creates a fresh $t^{{\text{th}}}$ sub-network dedicated to the current learning task, leveraging both the STFT features and the task metadata—such as task id—to discern previously unidentified acoustic configurations.
 What sets the $t^{{\text{th}}}$ sub-network apart is its ability to access the shared memory of its predecessors through lateral connections, integrating all previous knowledge.
-When it comes to inference application, the DOA-PNN model selects the appropriate $t^{{\text{th}}}$ sub-network based on the specific task at hand for accurate and reliable evaluation.
-This approach ensures that our model remains adaptable and quick to respond in diverse acoustic settings, showcasing the potential of progressive continual learning strategy in the real-world environments of DOA estimation.
-III Experiments
+When it comes to inference application, the CL-SSL model selects the appropriate $t^{{\text{th}}}$ sub-network based on the specific task at hand for accurate and reliable evaluation.
+This approach ensures that our model remains adaptable and quick to respond in diverse acoustic settings, showcasing the potential of continual learning strategy in the real-world environments of SSL.
+IV Experiments
 TABLE I: Parameter settings of simulated audio
@@
 [10,8,5], [5,4,3], [4,2,2] 0.2-0.8 0.2-0.6 [10,10,10]
-III-A Dataset
-We use the LibriSpeech [26] train-clean-100 subset, which comprises high-fidelity, single-channel speech recordings sampled at 16 kHz.
-It includes a diverse set of English speakers, selected for uniform audio segment lengths, ensuring homogeneity in data processing.
-From this dataset, we have curated a selection of 100 speakers out of the original 250.
-This selection balances gender representation with 47 male and 53 female speakers and encompasses a total of 1,511 audio files, amounting to approximately 4.96 hours of speech data.
-We enhance the original single-channel speech data by creating two-channel versions with gpuRIR [27], arranging two linear microphones at distances ranging from 0.05 to 0.09 meters.
-The gpuRIR is a free and open-source Python library for room impulse response (RIR) simulation using the image source method with GPU acceleration.
+IV-A Dataset
+We evaluated the proposed method on both simulated and real-world datasets, using two microphones to localize the direction of arrival within a 180-degree azimuth range.
+Simulated dataset: We create our simulated data based on the LibriSpeech [29], which comprises single-channel speech recordings sampled at 16 kHz.
+We have curated a selection of 100 speakers.
+This selection balances gender representation with 47 male and 53 female speakers and encompasses a total of 1,511 audio files (4.96 hours).
+We enhance the original single-channel speech data by creating two-channel versions with gpuRIR [30], arranging two linear microphones at distances ranging from 5 to 9 cm.
+The gpuRIR is a library for room impulse response (RIR) simulation using the image source method.
 The details of the simulated room environments are specified in Table I.
-To ensure sound sources are well-distributed across the room, we first simulate the microphone center anywhere inside the room.
+We first simulate the microphone center anywhere inside the room.
 Then we randomly choose a distance such that the source is inside the room to finally simulate the RIR.
 Based on the RIR, we generate the mixture data for each original single-channel speech data every 5 degrees from 0 to 180 degrees, which expands the audio files 37 times.
-III-B Metrics
-For evaluating our models, we utilize two primary metrics: mean absolute error (MAE) and accuracy (ACC).
+Real-world dataset: We also tested our proposed method on the LOCATA dataset [31] following the setting in [32].
+The room size was $7.1\times 9.8\times 3$ m, with a reverberation time of 0.55 s.
+We utilized microphones 6 and 9 from the DICIT array [33], spaced 8 cm apart.
+White, babble, and factory noises from the NOISEX-92 [34] dataset were employed as noise sources.
+The models, trained on a noise-augmented simulated dataset, were directly evaluated on the LOCATA dataset using the same microphone configuration.
+Various signal-to-noise ratio (SNR) levels were applied to showcase the CL-SSL performance on real-world datasets.
+IV-B Metrics
+We utilize two primary metrics for evaluating our models: mean absolute error (MAE) and accuracy (ACC).
+They are all averaged on all learned tasks.
 MAE measures the average magnitude of the errors in a set of predictions, without considering their direction.
 On the other hand, ACC is used to determine the percentage of predictions that are exactly correct or within a certain degree of tolerance.
-We assess accuracy at three levels of tolerance: Within a 5-degree, 10-degree, and 15-degree error margin.
-These tiers of tolerance allow us to understand not only the model’s precision but also its practical effectiveness.
+We assess accuracy at three tolerances: Within a 5-degree, 10-degree, and 15-degree error margin.
+These common tiers of tolerance allow us to understand the model’s precision and practical effectiveness.
 (a) ACC ±$\pm 5$
@@
 (c) ACC ±$\pm 15$
-Figure 2: Comparative inference performance of various DOA estimation methods across different microphone spacings, showing ACC within three tolerances.
-III-C Incremental learning setting
-In our incremental learning approach, we divide the dataset into five parts according to the varying distances between the microphones, with each partition representing a unique task.
-This allows us to treat the problem as a series of incremental learning tasks, each corresponding to a microphone spacing.
+Figure 2: Comparative performance in ACC (%) of various SSL methods after learning each microphone spacing for three tolerances levels. The tasks T1-T5 are microphones distance from 5 to 9 cm.
+IV-C Incremental learning setting
+In our incremental learning approach, the simulated dataset is divided into five parts based on microphone distances ranging from 5 to 9 cm.
+This setup frames the problem as a series of incremental learning tasks, each corresponding to a specific microphone spacing.
+For real-world data, noise levels are added using SNR values randomly selected from [-5dB, 0dB, 5dB, 10dB, Clean], representing different acoustic environments.
+The task order is randomly shuffled to better reflect practical scenarios.
 We compare our approach with several baseline methods:
   • Finetune: Starting with a model pre-trained on the initial task, we sequentially fine-tune it on each subsequent task.
-This approach adapts the model incrementally, but may be susceptible to forget previously learned information when moving to new tasks each time.
-  • Joint: A single model is trained on the entire dataset, encompassing all microphone spacings together. This method aims to learn a general representation that performs well across all the tasks.
-  • Ensemble: This involves training five separate models, each on a different subset of the data, with each model specializing in a particular microphone spacing.
-III-D Implementation details
-For training each task within our DOA-PNN framework, we consistently apply 100 epochs with an early stopping criterion set to activate after 20 epochs if there’s no improvement in performance.
-We employ the AdamW [28] optimizer, a variant of the Adam optimizer that includes weight decay for regularization, with a learning rate of 0.001.
-Additionally, we apply a learning rate scheduler, specifically StepLR, to adjust the learning rate during training.
-For the STFT computations, we set the frame size to 32 milliseconds and the hop size to 10 milliseconds.
-The frequency range for the STFT is confined between a minimum of 100 Hz and a maximum of 8000 Hz, using an n_fft value of 512, which determines the length of the windowed signal after padding with zeros.
-This STFT configuration is designed to capture a wide range of audio frequencies suitable for accurately determining the DOA.
-Gaussian function constant $\sigma$ we set as 8 in our experiments.
-TABLE II: Performance comparison across different DOA estimation methods, highlighting MAE, ACC within various degree tolerances, and model parameters.
-Methods MAE ACC(±$\pm 5$) ACC(±$\pm 10$) ACC(±$\pm 15$) Params
-Finetune 17.8 33.9 48.0 58.4 0.9M
-Joint 11.9 54.0 66.7 74.3 0.9M
-Ensemble 7.6 78.6 84.8 87.5 4.8M
-Proposed DOA-PNN 10.1 72.9 82.7 86.1 1.6M
-IV Results and Discussions
-We test the methods for DOA estimation across varying microphone spacings as Fig.
-2, we can observe how each method adapts to the acoustic challenge.
-The finetune method incrementally trained from 5cm to 9cm only performs well for the recent task 9 cm indicating a forgetting across previously learned configurations.
-On looking to the performance for joint training, we find that it is more resilience compared to finetune, yet it does not maintain high performance as spacing widens.
-It is possible that the single-model approach is not able to capture the nuances needed for localization at variable spacings.
-The ensemble method demonstrates a robust performance across different spacings as the upper bound.
-It benefits from training distinct models for each spacing scenario.
-Our proposed DOA-PNN shows a promising ability to sustain higher accuracy levels, particularly within the most challenging ±$\pm 5$-degree margin.
-It implies that our progressive learning strategy, which builds upon knowledge from previous configurations transfers well to new spacing scenarios.
-The proposed DOA-PNN offers a balance between accuracy and model complexity, suggesting it to be an efficient choice for real-world applications where microphone spacing can vary.
-On examining the comparative results of different methods for DOA estimation as given in Table II, we find each method has its strengths and limitations.
-The finetune approach, while the most straightforward in terms of adapting pre-existing models to new tasks, seems to lag in precision.
-It presents the highest MAE, suggesting that it might not preserve the nuances of DOA estimation in different microphone spacing as effectively as the other methods.
-The joint training method improves upon finetune significantly, indicating that learning from a collective dataset of all tasks offers a better understanding of the DOA estimation problem.
-However, this method still does not reach the top performance in accuracy, especially in ±$\pm 5$-degree range.
-Ensemble, which employs a separate model for each condition, stands out in accuracy.
-It’s clear that specializing in individual conditions allows for a finer understanding of each specific DOA estimation task, but this comes at the cost of a larger overall model size, as reflected by the parameter count.
-Our Proposed DOA-PNN method maintains a balance between precision and efficiency.
-It achieves comparable accuracy, nearly matching the best-performing method in the ±$\pm 15$-degree range and outperforming it within the ±$\pm 10$-degree margin.
-It achieves this while maintaining a moderate model size.
-This indicates that our progressive approach, which incorporates learning from successive tasks, effectively consolidates knowledge without increasing complexity.
+This approach adapts the model incrementally but may be susceptible to forgetting previously learned information when moving to new tasks each time.
+  • Joint: A single model is trained on the entire dataset, encompassing all conditions.
+This method aims to learn a general representation that performs well across all tasks.
+However, it can not handle large domain mismatches.
+  • Multi-Condition (MC): This involves training separate models for each condition on a different subset of the data, with each model specializing in a particular task.
+In incremental task setting, the performance can never surpass the MC as it already knows the whole target classes the first time and ensemble multiple models.
+  • Elastic Weight Consolidation (EWC) [35]: is a continual learning method that uses a quadratic penalty to restrict changes to parameters critical for previous tasks.
+These critical parameters are identified using an approximation based on the Fisher Information Matrix.
+  • Learning without Forgetting (LwF) [36]: is a method that ensures previously learned knowledge is retained while training on new tasks.
+It uses knowledge distillation, where the model’s predictions on old tasks are treated as “teacher” during the new task training.
+  • Memory Aware Synapses (MAS) [37]: is similar to EWC but identifies critical parameters by measuring their impact on output predictions.
+IV-D Implementation details
+In our CL-SSL framework, each task is trained for up to 100 epochs, with early stopping triggered after 20 epochs without performance improvement.
+We use the AdamW optimizer with weight decay, a learning rate of 0.001, and the StepLR scheduler for adjustments during training.
+For STFT computations, we use a frame size of 32 ms, a hop size of 10 ms, and a frequency range of 100 Hz to 8000 Hz, with n_fft set to 512.
+The Gaussian constant $\sigma$ is set to 8 in our experiments.
+V Results and Discussions
+V-A Results on simulated dataset
+We first test the methods for SSL across varying microphone spacing and report the results in Fig.
+2 and Table II.
+For the non-CL methods, we observe the finetune method incrementally trained only performs well for the first task indicating a forgetting across learned tasks.
+It presents the highest MAE, suggesting that it might not preserve the nuances of SSL in different microphone spacing as effectively as the other methods in Table II.
+Observing the performance of joint training, we find that it is more resilient compared to finetune, yet it does not maintain a high performance similar to the MC method.
+It is possible that the single-model approach is unable to capture the nuances needed for localization at variable spacing.
+The MC method shows consistent performance across different spacings due to the training of separate models for each task.
+However, it demands five times more memory and training time, making it less practical for real-world applications.
+As the number of tasks grows, the memory and time requirements increase linearly.
+For CL methods, all methods outperform finetuning, demonstrating their effectiveness in reducing forgetting.
+However, their performance remains inferior to that of joint training and MC.
+In addition, from Fig.
+2, it is observed that the performance of EWC and MAS drops dramatically when the number of learning tasks increases.
+This gap may result due to inappropriate regularization, which limits the model’s ability to learn new tasks while preserving prior knowledge.
+Among the reference CL methods, LwF performs better, likely because its use of the model’s predictions of old tasks enables knowledge sharing across tasks.
+Our proposed CL-SSL shows a promising ability to obtain higher accuracy levels, particularly within the most challenging ±$\pm 5$-degree margin as shown in Fig.
+2.
+It achieves comparable accuracy, nearly matching the best-performing method in the ±$\pm 15$-degree range in Table II.
+This indicates that our continual learning strategy effectively transfers knowledge from previous configurations to new spacing scenarios.
+By employing the layer scaling mechanism, we maintain a parameter count of just 1.6M—lower than LwF—while achieving significantly better performance.
+This highlights the proposed CL-SSL as a balanced solution between accuracy and model complexity, making it an efficient choice for SSL applications with varying microphone spacings.
+TABLE II: Performance comparison in ACC (%) across different methods on the simulated dataset, highlighting MAE. The performance is calculated after all tasks are learned.
+Methods MAE ACC(±$\pm 5^{\circ}$) ACC(±$\pm 10^{\circ}$) ACC(±$\pm 15^{\circ}$) Params
+Finetune 26.5 12.7 23.8 34.7 0.9M
+Joint 13.0 50.2 64.1 72.2 0.9M
+MC 7.6 77.6 85.6 88.4 4.5M
+EWC [35] 23.5 25.4 38.7 51.4 0.9M
+LwF [36] 22.1 35.5 47.4 58.9 1.8M
+MAS [37] 23.3 24.8 36.3 48.0 0.9M
+Proposed CL-SSL 10.1 71.5 83.0 87.0 1.6M
+V-B Results on real-world dataset
+As discussed, we further evaluate the proposed CL-SSL method on a noisy real-world dataset.
+Noise is added based on SNR values randomly selected from [-5 dB, 0 dB, 5 dB, 10 dB, Clean], representing different acoustic environments, with task order randomly shuffled for practicality.
+The results compared with other baselines are summarized in Table III.
+We observe better performance across all methods on the LOCATA dataset due to its more stable environment, including consistent room size, microphone center, and room impulse response.
+The finetune method remains the lower bound among non-CL methods, indicating that catastrophic forgetting negatively impacts performance across varied acoustic environments due to domain mismatches.
+Joint training achieves relatively better results on real-world data than on the simulated data (Table II), likely because optimizing for diverse acoustic environments is easier than for varying microphone configurations.
+The MC method, serving as the upper bound, performs well but at the cost of significantly higher memory usage.
+Among the CL baselines, EWC, LwF, and MAS show comparable performance and outperform the finetune method, suggesting that these approaches effectively mitigate forgetting while balancing adaptation to new tasks.
+Our proposed CL-SSL method demonstrates less forgetting compared to other CL methods and performs closely to the MC method.
+When compared to joint training, often regarded as the upper bound in many applications, CL-SSL achieves comparable results despite not having access to all tasks in advance.
+This suggests that our method has greater potential to perform well in unseen acoustic environments.
+Notably, in the most challenging ±$\pm 5$-degree scenario, CL-SSL even surpasses joint training.
+This advantage may stem from its ability to adaptively scale parameters and focus on critical features, enabling better fine-grained localization under strict error tolerances.
+TABLE III: Performance comparison in ACC (%) across different methods on the real-world dataset, highlighting MAE. The performance is calculated after all tasks are learned.
+Methods MAE ACC(±$\pm 5^{\circ}$) ACC(±$\pm 10^{\circ}$) ACC(±$\pm 15^{\circ}$) Params
+Finetune 17.2 43.0 58.6 67.7 0.9M
+Joint 5.2 68.4 78.9 86.8 0.9M
+MC 4.3 78.9 89.4 94.7 4.5M
+EWC [35] 9.3 47.3 68.4 81.5 0.9M
+LwF [36] 8.9 50.0 65.8 81.8 1.8M
+MAS [37] 10.8 48.8 67.1 77.4 0.9M
+Proposed CL-SSL 5.1 72.9 78.9 86.5 1.6M
+V-C Future Directions
 We foresee several avenues for future work.
-While we currently focus on two-microphone linear arrays to estimate the azimuth angle, we plan to expand our research to encompass more complex acoustic configurations and to address elevation angle estimation for a full spatial representation [29].
-Moreover, integrating noise robustness factors into DOA estimation is a critical next step, acknowledging that real-world applications often operate in acoustically challenging environments [30].
-These enhancements will aim to bolster the practicality of DOA-PNN, providing comprehensive solutions for 3D sound localization in noisy, real-life settings.
-V Conclusions
-In this study, we presented DOA-PNN, an approach that advances the direction of arrival estimation within continual learning.
-DOA-PNN synergizes task-specific network instantiation with progressive learning principles, addressing the diversity in microphone configurations with a novel scaling mechanism and lateral connections for knowledge retention.
-Demonstrated across simulated acoustic settings, DOA-PNN consistently achieved high accuracy, underscoring its capacity for real-world applicability in dynamic auditory scenes.
+While we currently focus on two-microphone linear arrays to estimate the azimuth angle, we plan to expand our research to encompass more complex acoustic configurations [38] and to address elevation angle estimation for a full spatial representation.
+VI Conclusions
+In this study, we presented CL-SSL, an approach that advances the direction of arrival estimation within continual learning.
+CL-SSL synergizes task-specific network instantiation with progressive learning principles, addressing the diversity in microphone configurations with a novel scaling mechanism and lateral connections for knowledge retention.
+Demonstrated across simulated acoustic settings, CL-SSL consistently achieved high accuracy, underscoring its capacity for real-world applicability in dynamic auditory scenes.
 The model’s efficiency and adaptability suggest its potential as a foundational framework for future research in sound localization and acoustic signal processing, encouraging the development of sophisticated, scalable audio analysis technologies.